A plane with a velocity 400 km/h, 10 degrees, West of South encounters a wind with a

velocity of 40 km/h, from 35 degrees East of South. What is the plane’s resultant

velocity?


1
Expert's answer
2022-04-29T14:17:06-0400
"\vec{v}=-400\sin10\degree \vec{i}-400\cos10\degree\vec{j}"

"\vec{u}=40\sin35\degree \vec{i}-40\cos35\degree\vec{j}"

"\vec{v}_{res}=\vec{v}+\vec{u}"

"=(-400\sin10\degree +40\sin35)\vec{i}"

"+(-400\cos10\degree-40\cos35\degree)\vec{j}"

"\approx-46.5162\vec{i}-426.6892\vec{j}"

"|\vec{v}_{res}|=\sqrt{(-46.5162)^2+(-426.6892)^2}"

"\approx429.2172(km/h)"

"\tan \theta=\dfrac{-426.6892}{-46.5162}"

"\theta=264\degree"

429.217 km/h, 6 degrees, West of South